OpenStudy (anonymous):

integrate e^x/(e^x-1) from -1 to 1

OpenStudy (anonymous):

forget that one. will not be finite

OpenStudy (anonymous):

put \[u=e^x,du=e^xdx\] get \[\int\frac{du}{u}=\ln(u)=\ln(e^x-1)\]

OpenStudy (anonymous):

this is an improper integral, because the path includes 0 and your integrand is not defined at 0. so you have to compute \[\int_r^1\frac{e^x}{e^x-1}dx\] and then take the limit as r goes to 0

OpenStudy (anonymous):

but \[\lim_{r\rightarrow 0^+} \ln(e^r-1)\] does not exist , so no limit and hence no integral.